OpenStudy (anonymous):

Write the standard form of the equation of each line passing through P with slope m. P=(0,6); m=1/1

OpenStudy (marissalovescats):

Use the equation y-y1=m(x-x1) and then put that into y=mx+b form :)

OpenStudy (marissalovescats):

Let me know if you need any help or guidance. I'm here.

OpenStudy (marissalovescats):